OpenStudy (anonymous):

Which of these are rational numbers? a. 3.75 b. -132 c. .140625 d. .04

OpenStudy (jamesj):

What's the definition of a rational number?

OpenStudy (anonymous):

You tell me ?

myininaya (myininaya):

\[\frac{integer_1}{integer_2}, integer_2 \neq 0\]

OpenStudy (jamesj):

Now, given that is your first number 3.75 rational?

OpenStudy (anonymous):

ya ?

OpenStudy (jamesj):

A rational number is a number that can be written as a ratio or fraction of two integers. For example. All integers themselves are rational numbers. 1 = 1/1, a ratio of two numbers. 2 = 2/1, 3 = 3/1. -1 = -1/1, -2 = -2/1, etc. All finite decimal expansions are rational numbers. E.g. 1.5 = 15/10 2.34 = 234/100 There are other rational numbers as well, such as 1/7. ===== Now, given that is your first number 3.75 rational?

OpenStudy (jamesj):

Yes. 3.75 = 375/100. That's a fraction of two integers. Therefore 3.75 is a rational number.

OpenStudy (jamesj):

What about -132 ?

OpenStudy (anonymous):

no. theres only one rational number in that group of numbers

OpenStudy (jamesj):

That's not true. Didn't we just say that every integer is itself a rational number?

OpenStudy (jamesj):

-132 = -132/1 hence -132 is also a rational number.

myininaya (myininaya):

James is right there is more than one rational number here.

OpenStudy (jamesj):

What about part c. 0.140625 Rational or not rational?

OpenStudy (jamesj):

And you've left. Ok. Good luck then.